# Cybersecurity in Critical Infrastructures: A Post-Quantum Cryptography Perspective

### v1

@@ IV.6 Multiparty Computation protocol and Graph-based cryptography
In the PQC standardization processes around the word only South Korea is still considering protocols based on these problems as we will see in the next section [103].
IV.6.1 Multiparty Computation protocol
@@
Multiparty Computation Protocol (MPC) cryptography is used in scenarios where several parties $P_{i}$ want to make some data available keeping its confidentiality. As an example, we could imagine two different countries which wants to keep the trajectory of its spy satellite secret but they want to be sure that they have different trajectory in order to avoid collisions. Therefore, each party has a secret $x_{i}$, the trajectory of their satellite, which has to remain secret but they want to share some confidential information making sure they will not collide. This is known as the MPC problem. In this sense, the MPC paradigm relates with the interest of allowing some party to do some computations to extract some conclusions using some protected data without actually having access to the raw data. The zero-knowledge proof allows the entities to convince the other ones about something without making the data public. The first MPC scheme was proposed by Y. Ishai, E. Kushilevitz, R. Ostrovsky and A. Sahai in 2007 [125], known as MPC-in-the-head paradigm. The first application proof of this paradigm was presented by Giacomelli et al. at 2016 in [126] and protocols based on this paradigm can be used to generate signatures. Picnic [127] is a PQC signature scheme that has been proposed for the NIST standardization process, but that was rejected in the third round [101]. However, South Korea keeps a PQC algortihm based on MPC-in-the-head paradigm in its standardization process called AIMer [128].
IV.6.2 Graph-based Cryptography
@@
Graph-based cryptography refers to Perfect Code Cryptosystems (PCC) proposed by Koblitz et al. in 1993 [129]. The study of Perfect Codes (PC) emerged in the field of information theory, since PCs over graphs corresponds to PCs over structured alphabets [130]. Despite of having its origin in the study of codes, we separate this cryptography from code-based cryptography explained in section IV.3 since the hard problem to be solved in order to break each of the cryptosystems is different. A graph is a mathematical object formed by a set of edges and vertices, where each vertex is connected to other vertices through edges. A Perfect Code (PC) in a graph is defined as the set of vertices $A$ such that every vertex not included in $A$ is connected to only one element in $A$. A graph is defined as $G={V,E}$, where $V$ is a set of vertices and $E$ is a set of edges. Therefore, $A\subseteq V$ and for every $v\in V$, $N[v]$ contains only one element of $A$, where the set $N[v]$ is formed by the vertices connected by an edge to $v$. The security of this cryptosystem relies on the hardness of knowing if a graph has PCs or not, which is a hard problem by itself (NP-complete problem) [130], as well as on the hardness of finding the vertices which form the PC, which is conjectured to be a hard problem (NP-complete problem). IPCC [131] is a graph-based post-quantum cryptography protocol proposed to be standardized in the South-Korean standardization process, which improves the original ideas proposed by Koblitz.
IV.7 Performance of PQC algorithms
@@
The NIST has defined 5 security levels to compare the performance of the different PQC algorithms when the security they provide is the same [132]. Some metrics on the performance of the different PQC algorithms proposed to be standardized to achieve first level of security are presented in the appendix A in the tables 3 (Lattice-based), 4 (Hash-based), 5 (Code-based), 6 (Multivariate cryptography), 7 (Isogeny-based) and 8 (MPC and Graph-based). In general terms their performance in OT communications can be briefly discussed:
@@
  • Lattice-based (Table 3): Lattice-based cryptography requires relatively small public and private keys as well as ciphertexts while showing a lower computational cost compared to other PQC families [133]. Hence, it is the most promising candidate to be implementable in ICS/CI scenarios.
@@
  • Hash-based (Table 4): There is only one hash-based PQC protocol, Sphincs+, proposed along the world, and its problems are the large ciphertext and the lengthy signing time [134], which could make it not suitable for ICS/CI.
@@
  • Code-based (Table 5): Their public and private keys’ length are longer than for other PQC families implying that more memory is required. Also, the fact that deciphering the ciphertext requires a high computational cost makes code-based not to be the best option to be deployed in an ICS/CI environment a priori [101].
@@
  • Multivariate-based (Table 6): Multivariate-based cryptosystems require large keys while their ciphertext lengths are the same as for other PQC families, or even shorter. Nonetheless, due to the high computational and memory resources requirements [133] multivariate cryptography is not the best option to be used in CI.
@@
  • Isogeny-based (Table 7): Isogeny based cryptography presents keys of moderate length, however, FIBS 7 has a large signature, making it unappealing to be implemented in CI. Although isogeny-based protocols have small key sizes, the computational cost is a great disadvantage due to the low computational requirements in ICS/CI.
@@
  • Graph-based and multiparty computation cryptography (Table 8): These are the least used PQC algorithms along the world. They are not receiving a lot of attention and, thus, their security against classical and quantum attacks are not as studied as for other PQC families. Due to the nature of ICS/CI is not in principle recommended to use protocols based on these hard problems for such contexts.
@@
The result of this high level comparison among PQC algorithms implies that, a priori, the most suitable protocols for its integration in ICS/CI networks are those that belong to the lattice-based cryptography family. However, and as discussed before, this conclusion is rather speculative and, thus, an actual comparison in the conditions of those scenarios should be done for obtaining accurate results and conclusions because it is probably that any of them satisfies all industrial infrastructure communication’s constrains.
@@
SMAUG [85]
LWE Signature Crystal-Dilithium [87]
@@
Crystal-Dilithium [87] - Crystal-Dilithium [87] Crystal-Dilithium [87]
GCKSign [90]
@@
Small-Field Rainbow [112] - Rainbow [112] - Rainbow [112] - -
@@
Perfect Code - - - - - - IPCC [131]
Table 1: The data have been obtained from ETSI (European Telecommunications Standards Institute) in the case of Europe [136]; from CACR (Chinese Association for Cryptologic Research) in the case of china [137]; from NIST (National Institute of Standards and Technology) in the case of United States of America [101]; from BSI (Bundesamt für Sicherheit in der Informationstechnik) in the case of Germany from CRYPTREC (Cryptography Research and Evaluation Committees) in the case of Japan [138]; from ANSSI (Agence Nationale de la Sécurité des Systèmes d’Information) in France [139] and from KpqC (Korean post-quantum Cryptography) in the case of South Korea [103]
@@
Table 2: Global distribution of the public investment in quantum technologies in 2023 [140] and in 2022 [141]. The total investment was around 27 billion euros in 2022 and increased a $33\%$ in 2023, to 36 billion euros.
IV.8 PQC standardization processes around the world
@@
Despite of the global concern regarding the cybersecurity threat posed by the possibility of constructing fault-tolerant quantum computers, not every country/entity has started a standardization process by their own, or if they has started they do not publish this information. In this context, the United States is the pioneer by means of the NIST post-quantum cryptography standardization process, which started in 2017 and is yet unfinished (at the moment they are conducting the fourth round [24]). Even if it stands as the largest standardization process, not all countries will adopt NIST recommendation and standardized PQC algorithms as a result of economical and political differences. Countries not aligned with the North Atlantic Treaty Organization (NATO) and the USA foreign policy are developing their own algorithms such as, for example, China. However, even EU countries that belong to such alliances are developing PQC protocols and requirements by their own, e.g. Germany and France. There are other countries which are investing in quantum technologies but they do not have published a PQC standardisation process. However, by comparing the table 1 with 2 it can be inferred that it is quite probable that they are making their own efforts to be quantum secure.
@@
The main issue with the adoption of PQC schemes for cryptography tasks resides on the youth of most of the proposals. More concretely, there is not a wide experience in the integration of the schemes in real systems implying that the security of the experimental implementation of PQC schemes is still a question to be deeply explored. Due to the criticality of cybersecurity regarding the security of national secrets, most governments are still skeptical of completely relying on those novel methods. This is also a result of the fact that many of the cryptography protocols proposed in the past were proved to be insecure years later of their proposal. This is somehow result of the fact that some of the hard problems in which cryptography is based on rely on mathematical assumptions regarding their hardness, which is a evolving science. Obviously, this problem is exacerbated when PQC security proofs are considered, mainly because the class of problems that are solvable by quantum computers and its relationship to other complexity classes is still a question under research. Also, even with formal proofs of the hardness, the actual implementations of the protocols are not guaranteed to be safe. Therefore, due to the early stage of PQC proposals, many countries are exploring different possibilities as a function of their own analyses and interests rather than relying on the recommendations by a single entity such as the NIST. In Table 1 and Table 2, we present the global efforts regarding PQC integration in their communications and public investments, respectively, in quantum technologies as a way of showing a picture of the state of affairs at the time of writing.
@@ V Post-Quantum cryptography in Critical Infrastructure
Providing cybersecurity resources to CIs is an indispensable task due to the fact that their constituting elements are interconnected among them and with other CI industries, implying that a weakness in any point of the network could produce a cascading effect that would result in a large economical, social and human cost, as we have explained in the section II. An adversary with the computational power of a quantum computer could take advantage of weakness in a concrete part of CI’s communications to launch a fatal quantum attack that could potentially affect several urban centers, industrial and state infrastructures. Therefore, integrating classical cryptographic methods in CI networks will provide security nowadays, but since the life-time span of OT devices is longer than the time scale estimated for the construction of the first fault tolerant quantum computer [^6] Note that it is projected that, at the current pace, IBM quantum processors could crack RSA by 2040 [142]., such solution can be deemed as a patch in the goal of protecting CI networks. Consequently, PQC solutions for CI are a crucial necessity and, hence, it is necessary target PQC from the point of view of CIs, fulfilling the required stringent communications requirements with the low computational resources/legacy devices and testing the proposed solutions in real environments.
@@
The most important characteristic a PQC algorithm should present in order to be suitable for its implementation in CI, in combination with security, is a low computational time (as this relates to the latency added to the system). This comes from the fact that some controlling operations in OT require latencies of the order of milliseconds, as in IEC-62443 [34], failing to satisfy such latency constraints may cause failures on the system. However, since each of the discussed PQC protocols have not been tested under the same conditions (processor, benchmark, security level …), performing a high-level comparison by means of the provided latencies would be inaccurate. Thus, the conclusions would not be relevant for the application of the protocols in ICS/CI. This also comes in hand with the fact that since such implementations have not been realized from the point of view of ICS/CI, i.e. trying to fulfill the stringent conditions imposed by such systems, the obtained conclusions would only be partially true. Despite of the lack of such benchmark, it is possible to somehow bound the performance of PQC families in order to select which of them could be a potentially good option to be deployed in an ICS system. We have done this by using the tables of the appendix A in the discussion in IV.7, where we compare the keys and ciphertext length of each PQC algorithm and their computational cost. However, the lack of actual fair comparative metrics for PQC in industrial and critical infrastructure networks urges for performing such comparison in the same conditions and from the point of view of the necessities of such scenarios. In this way, the selection of PQC protocols for deployment in ICS/CI will be actually possible, feature that has recently been pointed out by the CISA, NSA and NIST to be of critical importance [33].

### v2

@@ IV-F Multiparty Computation protocol and Graph-based cryptography
In the PQC standardization processes around the word only South Korea is still considering a protocol based on one of these problems as we will see in the next section [136].
IV-F1 Multiparty Computation protocol
@@
Multiparty Computation Protocol (MPC) cryptography is used in scenarios where several parties $P_{i}$ want to make some data available keeping its confidentiality. As an example, we could imagine two different countries which wants to keep the trajectory of its spy satellite secret but they want to be sure that they have different trajectory in order to avoid collisions. Therefore, each party has a secret $x_{i}$, the trajectory of their satellite, which has to remain secret but they want to share some confidential information making sure they will not collide. This is known as the MPC problem. In this sense, the MPC paradigm relates with the interest of allowing some party to do some computations to extract some conclusions using some protected data without actually having access to the raw data. The zero-knowledge proof allows the entities to convince the other ones about something without making the data public. The first MPC scheme was proposed by Y. Ishai, E. Kushilevitz, R. Ostrovsky and A. Sahai in 2007 [165], known as MPC-in-the-head paradigm. The first application proof of this paradigm was presented by Giacomelli et al. at 2016 in [166] and protocols based on this paradigm can be used to generate signatures. Picnic [167] is a PQC signature scheme that has been proposed for the NIST standardization process, but that was rejected in the third round [134]. However, AIMer[168] PQC algorithm based on MPC-in-the-head paradigm had been selected for the second round in the South Korea PQC standardization process.
IV-F2 Graph-based Cryptography
@@
Graph-based cryptography refers to Perfect Code Cryptosystems (PCC) proposed by Koblitz et al. in 1993 [169]. The study of Perfect Codes (PC) emerged in the field of information theory, since PCs over graphs corresponds to PCs over structured alphabets [170]. Despite of having its origin in the study of codes, we separate this cryptography from code-based cryptography explained in Section IV-C since the hard problem to be solved in order to break each of the cryptosystems is different. A graph is a mathematical object formed by a set of edges and vertices, where each vertex is connected to other vertices through edges. A Perfect Code (PC) in a graph is defined as the set of vertices $A$ such that every vertex not included in $A$ is connected to only one element in $A$. A graph is defined as $G={V,E}$, where $V$ is a set of vertices and $E$ is a set of edges. Therefore, $A\subseteq V$ and for every $v\in V$, $N[v]$ contains only one element of $A$, where the set $N[v]$ is formed by the vertices connected by an edge to $v$. The security of this cryptosystem relies on the hardness of knowing if a graph has PCs or not, which is a hard problem by itself (NP-complete problem) [170], as well as on the hardness of finding the vertices which form the PC, which is conjectured to be a hard problem (NP-complete problem). IPCC [171] is a graph-based post-quantum cryptography protocol proposed to be standardized in the South-Korean standardization process, which improves the original ideas proposed by Koblitz.
IV-G Performance of PQC algorithms
Lattice-based Public Key Private key Ct/Signature
NTRU [172, 134] 930 1234 930
NTRU-HRSS [173, 134] 1 138 1 450 1 138
NTRU-Prime [173, 174, 134] 994 15 158 897
NTRU+ [94, 175] 864 1 728 864
Falcon citeFalconPerf,NIST3 897 7 553 666
FatSeal [96] 2 321 385 2 048
Peregrine [97, 175] 897 7 553 666
SOLMAE [98, 175] 897 7 553 666
CRYSTALS-Kyber [176, 134] 800 1 632 768
Saber [173, 134] 672 832 736
FrodoKEM [100, 134] 9 619 19 888 9 720
LAC PKE [101] 544 1 056 704
Aigis-Enc [102, 177] 672 1 568 672
AKCN-MLWE (AES256) [103] 10 560 10 560 8 610
TALE [104, 178] 736 1 504 704
AKCN-E8 (AES256) [105] 928 928 896
SCloud (AES256) [106] 98 800 19 800 82 300
TiGER [175, 108] 480 177 640
SMAUG [175, 107] 174 672 768
CRYSTALS-Dilithium(AES192) [134] 1 312 2 528 2 420
Aigis-Sig [110] 672 1 568 672
Mulan [111] 1 312 2 528 2 420
GCKSign [179, 175] 1 760 288 1 952
NCC-Sign [112, 175] 1 440 2 400 2 529
HAETAE [113, 175] 992 1 376 1 463
TABLE I: Performance of Lattice-based PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [180]
Hash-based Public Key Private key Ct/signature
SPHINCS+ [82, 176, 134] 32 64 7 856
TABLE II: Performance of Isogeny-based PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [180]
Code-based Public Key Private key Ct/Signature
BIKE [134] 1 540 280 1 572
PALOMA [135, 175] 319 488 7 808 32
Cl. McEllice [134] 261 129 6 492 128
HQC [134] 2 249 40 4 481
Piglet-1 [181] 1 212 32 1 801
ROLLO-I [182, 175] 1 240 120 620
REDOG [139, 175] 14 250 1 450 830
Enhanced PQsigRm [183, 175] 2 000 000 1 032
TABLE III: Performance of Code-based PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [180]
Multivariate Public Key Private key Ct/Signature
GeMSS [176, 134] 352 168 16 33
Rainbow [146, 134] 157 800 611 300 164
MQ-Sign [149, 175] 328 441 15 561 134
TABLE IV: Performance of Multivariate PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [180]
Isogeny-based Public Key Private key Ct/Signature
SIKE [134] 330 374 346
FIBS [160] 32 64 17 088
TABLE V: Performance of Isogeny-based PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [180]
Others Public Key Private key Ct/Signature
IPCC (AES80)[171] 4 800 400 92 000
AIMer [168] 32 16 5 904
TABLE VI: Performance of Isogeny-based PQC protocols. Ciphertext (Ct) and keys length are expressed in Bytes. All values are taken from the level I security defined by NIST: ”Any attack that breaks the relevant security definition must require computational resources comparable to or greater than those required for key search on a block cipher with a 128-bit key (e.g. AES128)” [180]
@@ IV-G Performance of PQC algorithms
The NIST has defined 5 security levels to compare the performance of the different PQC algorithms when the security they provide is the same [180]. Some metrics on the performance of the different PQC algorithms proposed to be standardized to achieve first level of security are presented in: Table I (Lattice-based), Table II (Hash-based), Table III (Code-based), Table IV (Multivariate cryptography), Table V (Isogeny-based) and Table VI (MPC and Graph-based). In general terms their performance in OT communications can be briefly discussed:
PQC family Key lengths Ct/Signature length Number of PQC algorithms Time* Appropriate for CI
Hash-based A D C D D
Lattice-based B B A A A
Code-based C B B D D
Multivariate D A B D C
Isogeny-based A B C C B
Multiparty A C D D D
Graph-based C D D C D
TABLE VII: Adequacy of PQC protocols to CI in terms of their key and ciphertext/Signature lengths, the number of PQC algorithms proposed to be standardized and the average computational time. The letters and the colors define the suitability of the PQC family, being the most suitable (A/green), moderately suitable (B/yellow), hardly suitable (C/orange) and not suitable (D/red).
@@
  • Lattice-based (Table I): Lattice-based cryptography requires relatively small public and private keys as well as ciphertexts while showing a lower computational cost compared to other PQC families [184]. Hence, it is the most promising candidate to be implementable in ICS/CI scenarios.
@@
  • Hash-based (Table II): There is only one hash-based PQC protocol, Sphincs+, proposed along the world, and its problems are the large ciphertext and the lengthy signing time [176], which could make it not suitable for ICS/CI.
@@
  • Code-based (Table III): Their public and private keys’ length are longer than for other PQC families implying that more memory is required. Also, the fact that deciphering the ciphertext requires a high computational cost makes code-based not to be the best option to be deployed in an ICS/CI environment a priori [134].
@@
  • Multivariate-based (Table IV): Multivariate-based cryptosystems require large keys while their ciphertext lengths are the same as for other PQC families, or even shorter. Nonetheless, due to the high computational and memory resources requirements [184] multivariate cryptography is not the best option to be used in CI.
@@
  • Isogeny-based (Table V): Isogeny based cryptography presents keys of moderate length, however, FIBS has a large signature, making it unappealing to be implemented in CI. Although isogeny-based protocols have small key sizes, the computational cost is a great disadvantage due to the low computational requirements in ICS/CI.
@@
  • Graph-based and multiparty computation cryptography (Table VI): These are the least used PQC algorithms along the world. They are not receiving a lot of attention and, thus, their security against classical and quantum attacks are not as studied as for other PQC families. Due to the nature of ICS/CI is not in principle recommended to use protocols based on these hard problems for such contexts.
@@
The result of this high level comparison among PQC algorithms implies that, a priori, the most suitable protocols for its integration in ICS/CI networks are those that belong to the lattice-based cryptography family. It is important to note that during the writing of this perspective, a new preprint was posted proposing a quantum algorithm that solves LWE problem with a polynomial complexity [185]. While the method proposed does not break NIST PQC candidates, developments in this field may lead to such outcome. Thus, even if out speculative comparison leans towards them, this advances in quantum algorithms should be taken into account. The comparison is summarized in Table VII. However, and as discussed before, this conclusion is rather speculative and, thus, an actual comparison in the conditions of those scenarios should be done for obtaining accurate results and conclusions because it is probably that any of them satisfies all industrial infrastructure communication’s constrains.
@@
LWE Signature CRYSTALS-Dilithium [109]
@@
CRYSTALS-Dilithium [109] - CRYSTALS-Dilithium [109] CRYSTALS-Dilithium [109]
@@
Small-Field - - - - - - -
@@
Perfect Code - - - - - - -
TABLE VIII: The data have been obtained from ETSI (European Telecommunications Standards Institute) in the case of Europe [186]; from CACR (Chinese Association for Cryptologic Research) in the case of china [187]; from NIST (National Institute of Standards and Technology) in the case of United States of America [134]; from BSI (Bundesamt für Sicherheit in der Informationstechnik) in the case of Germany[46]; from CRYPTREC (Cryptography Research and Evaluation Committees) in the case of Japan [188]; from ANSSI (Agence Nationale de la Sécurité des Systèmes d’Information) in France [45] and from KpqC (Korean post-quantum Cryptography) in the case of South Korea [136]. A discussion comparing their performances is given in Section IV-G, while a set of tables presnting the performance is given in Section IV.
@@
TABLE IX: Global distribution of the public investment in quantum technologies in 2023 [189] and in 2022 [190]. The total investment was around 27 billion euros in 2022 and increased a $33\%$ in 2023, to 36 billion euros.
IV-H PQC standardization processes around the world
@@
Despite of the global concern regarding the cybersecurity threat posed by the possibility of constructing fault-tolerant quantum computers, not every country/entity has started a standardization process by their own, or if they has started they do not publish this information. In this context, the United States is the pioneer by means of the NIST post-quantum cryptography standardization process, which started in 2017 and is yet unfinished (at the moment they are conducting the fourth round [23]). Even if it stands as the largest standardization process, not all countries will adopt NIST recommendation and standardized PQC algorithms as a result of economical and political differences. Countries not aligned with the North Atlantic Treaty Organization (NATO) and the USA foreign policy are developing their own algorithms such as, for example, China. However, even EU countries that belong to such alliances are developing PQC protocols and requirements by their own, e.g. Germany and France. There are other countries which are investing in quantum technologies but they do not have published a PQC standardisation process. However, by comparing the Table VIII with Table IX it can be inferred that it is quite probable that they are making their own efforts to be quantum secure.
@@
The main issue with the adoption of PQC schemes for cryptography tasks resides on the youth of most of the proposals. More concretely, there is not a wide experience in the integration of the schemes in real systems implying that the security of the experimental implementation of PQC schemes is still a question to be deeply explored. Due to the criticality of cybersecurity regarding the security of national secrets, most governments are still skeptical of completely relying on those novel methods. This is also a result of the fact that many of the cryptography protocols proposed in the past were proved to be insecure years later of their proposal. This is somehow result of the fact that some of the hard problems in which cryptography is based on rely on mathematical assumptions regarding their hardness, which is a evolving science. Obviously, this problem is exacerbated when PQC security proofs are considered, mainly because the class of problems that are solvable by quantum computers and its relationship to other complexity classes is still a question under research. Also, even with formal proofs of the hardness, the actual implementations of the protocols are not guaranteed to be safe. Therefore, due to the early stage of PQC proposals, many countries are exploring different possibilities as a function of their own analyses and interests rather than relying on the recommendations by a single entity such as the NIST. In Table VIII and Table IX, we present the global efforts regarding PQC integration in their communications and public investments, respectively, in quantum technologies as a way of showing a picture of the state of affairs at the time of writing.
Recently, South Korea has finished round 1 of their PQC Standardisation process [136]. After the whole process that has taken almost one year, they have decided to discard ROLLO-1 [182], Enhanced pqsigRm [183], SMAUG [107], GCKSign [179], TiGER [108], Peregrine [97], SOLMAE [98], FIBS [160] and IPCC [171]. On the contrary, the signature algorithms AIMer [168] as a MPC signature; HAETAE [113] and NCC-Sign [112] as lattice-based signatures; and MQ-Sign [149] as Multivariate signature have been maintained as candidates for the second round. Regarding PKE/KEM algorithms they have selected NTRU+ and a merge of SMAUG [107] and TiGER [108] as lattice-based; and PALOMA [135] and REDOG [139] as code-based for the second round.
@@ V Post-Quantum cryptography in Critical Infrastructure
Figure 7: A comparison of the lifespan of secrets in IT environments and the lifespan of devices in OT environments reveals significant differences. In IT communications, security must withstand quantum attacks several years prior to the advent of fault-tolerant quantum computers. This preemptive security measure is essential due to the risk of intercepting and storing communications today for decryption at a later time, known as ’Harvest now, decrypt later’ attack. Conversely, in OT communications, this particular issue is less pressing. However, cybersecurity in OT environments must still be quantum resilient, primarily because OT devices have long lifespans.
@@
Providing cybersecurity resources to CIs is an indispensable task due to the fact that their constituting elements are interconnected among them and with other CI industries, implying that a weakness in any point of the network could produce a cascading effect that would result in a large economical, social and human cost, as we have explained in the Section II. An adversary with the computational power of a quantum computer could take advantage of weakness in a concrete part of CI’s communications to launch a fatal quantum attack that could potentially affect several urban centers, industrial and state infrastructures. Therefore, integrating classical cryptographic methods in CI networks will provide security nowadays, but since the life-time span of OT devices is longer than the time scale estimated for the construction of the first fault tolerant quantum computer [^9] Note that it is projected that, at the current pace, IBM quantum processors could crack RSA by 2040 [191], such solution can be deemed as a patch in the goal of protecting CI networks. Fig. 7, shows the essential difference of the problems of securing IT and OT environments. The main point in trying to accelerate the integration of PQC in IT networks is the “harvest now, decrypt later” paradigm, i.e. possible attackers store encrypted data for decrypting it once a quantum processor is available. Therefore, a fast integration is required for IT from the point of view of confidentiality, since the attackers are interested in the actual content of the encrypted data. On the other hand, in OT environments a possible attacker is not interested in reading the content of the information being communicated, but it is interested in being able to violate the system, i.e. to break data authenticity and integrity to attack the critical infrastructure. Thus, the “harvest now, decrypt later” paradigm is not very relevant in this scenario. The big problem lays in the lifespan of the devices. As discussed before, industrial networks are assumed to last many decades and consist of legacy equipment, so integrating a secure solution against classical attacks may not be useful once quantum attacks can be realized, putting the entire system into a high degree of vulnerability. Protecting CI networks with quantum resilient solutions should aim to make them secure over the whole lifespan of the system. Consequently, PQC solutions for CI are a crucial necessity and, hence, it is necessary to target PQC from the point of view of CIs, fulfilling the required stringent communications requirements with the low computational resources/legacy devices and testing the proposed solutions in real environments.
One of the main issues with implementing PQC in communication networks, both in OT and IT, is the increased duration of the handshake between parties. This is primarily due to the larger key length of PQC cryptosystems compared to traditional protocols. The handshake between parties typically occurs by means of the Transport Layer Security (TLS) protocol. There are experiments documented in the literature that compare the handshake times between PQC cryptosystems and classical ones, as noted in [192]. The latency increase in IT communications poses a significant challenge, especially when there is a high volume of communication relying on this protocol. However, in the case of OT communications, an increase in latency is not merely an optimization concern; it could potentially result in fatal errors. Therefore, the most important characteristic a PQC algorithm should present in order to be suitable for its implementation in CI, in combination with security (data integrity and authenticity), is a low computational time (as this relates to the latency added to the system). This comes from the fact that some controlling operations in OT require latencies of the order of milliseconds, as in IEC-62443 [47], failing to satisfy such latency constraints may cause failures on the system.
Another important concern in OT cybersecurity are the Side-Channel Attacks. A side-channel attack is a type of security breach that involves analyzing patterns of information leakage from a system to gain unauthorized access to sensitive data. Instead of directly attacking the cryptographic algorithm itself, side-channel attacks exploit unintended side channels such as power consumption, electromagnetic emissions, acoustic emanation, or timing information to infer secret information, such as encryption keys. A comprehensive analysis and definition of each type of Side-Channel Attack (SCA) is provided in [193].
By observing these side channels, attackers can deduce valuable information about the internal operations of a system and potentially compromise its security. While side-channel attacks are a crucial consideration in IT cybersecurity, they are less of a concern in Operational Technology (OT) environments. This is because conducting a side-channel attack in OT environments typically requires physical access to the devices or prior infection with malware. However, nowadays physical access is not always necessary to conduct a side-channel attack. Some attacks can be executed remotely or with limited physical proximity to the target system, assuming that the device has been priory infected by a malware which send enough information to the attacker for doing a SCA. However, the feasibility and effectiveness of a side-channel attack may vary depending on the specific type of attack and the level of access to the target device. Although physical access may facilitate certain types of side-channel attacks on ICS, it is not always a strict requirement for successful attacks. For example, in [194] they perform a SCA to a infected PLC knowing its cache behaviour. This paper contributes to a better understanding of the risks posed by SCA in industrial control environments and emphasize the need for robust countermeasures to protect CI against this kind of attacks. Another important issue related to SCA is error and fault detection [195]. An error in the PQC algorithm could leak enough information to enable a SCA attack. For instance, some research focuses on enhancing the NTT [121, 196], which can reduce the computation time of lattice-based algorithms [197], but it may also leak information. A good review of NTT and its applications in PQC is given in [120]. Error and fault detection present a significant challenge not only in lattice-based cryptography but also in other PQC families, such as hash-based algorithms [198, 199].
@@
Many research efforts aiming the implementation of PQC algorithms in HW are being conducted by the community, as pointed out in Section IV-G, and examples of these implementations are given in the subsection ’PQC protocols’ for each PQC family (Sections IV-A2IV-B2IV-C2IV-D2IV-E2 and  IV-F. However, since each of the discussed PQC protocols has not been tested under the same conditions (processor, benchmark, security level …), performing a high-level comparison by means of the provided latencies would be inaccurate. Thus, the conclusions would not be relevant for the application of the protocols in ICS/CI. This also comes in hand with the fact that since such implementations have not been realized from the point of view of ICS/CI, i.e. trying to fulfill the stringent conditions imposed by such systems, the obtained conclusions would only be partially true. Recently, a study on implementing CRYSTALS-Kyber and CRYSTALS-Dilithium in IoT environments, meaning environments with limited computational resources, has been published [200]. These PQC algorithms were selected in the NIST standardization process [23]. The study pointed out the challenges of implementing PQC algorithms in IoT environments and presented an efficient and innovative lattice-based cryptography processor to make them suitable for IoT environments. However, more studies and implementations of PQC algorithms in IoT infrastructures are needed, particularly concerning SCA and the potential delay they could introduce to communication processes. Despite of the lack of such benchmark, it is possible to somehow bound the performance of PQC families in order to select which of them could be a potentially good option to be deployed in an ICS system. We have done this by using the tables in section IV-G, where we compare the keys and ciphertext length of each PQC algorithm and their computational cost, characteristics that, in the end, are related with the latency introduced to the communications and the requirements of the hardware used in the network. However, the lack of actual fair comparative metrics for PQC in industrial and critical infrastructure networks urges for performing such comparison in the same conditions and from the point of view of the necessities of such scenarios. In this way, the selection of PQC protocols for deployment in ICS/CI will be actually possible, feature that has recently been pointed out by the CISA, NSA and NIST to be of critical importance [32].
Another challenge in implementing any new cryptosystem into a standardized communication protocol is that it may require changes across all systems [33]. All systems must adopt the same cryptosystem for a successful handshake initiation, necessitating a migration of all systems from the classical TLS protocol to a quantum-secure protocol. However, not all systems are prepared to incorporate PQC in their current state. Some are old and lack the resources, while others may not be designed to accommodate this type of cryptography, even in IT communication[201]. In the context of OT communications, the fact that all devices have to adopt the same PQC protocol is less problematic due to the confined nature of communications within an industry. Communication between devices typically occurs only within the same industrial setting, necessitating standardization only within that specific industry. For instance, a PLC primarily communicates with other devices within the manufacturing control layer or the Area Supervisory Control layer (refer to Fig. 2), rather than with devices outside of this network. In order to communicate with elements outside this network, the protocols used are IT standards and, thus, do not concern the discussions presented here.